Use linear combinations/elimination to solve the following system. (It must be rearranged first)
4x=5y-14
3y-8x=-14

Answer by mananth(16946)

4 x -5 y = -14 .............1
Total value
-8 x + 3 y = -14 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 5
12 x -15 y = -42
-40 x + 15 y = -70
Add the two equations
-28 x = -112
/ -28
x = 4
plug value of x in (1)
4 x -5 y = -14
16 -5 y = -14
-5 y = -14 -16
-5 y = -30
y = 6